Margo borrows $1200, agreeing to pay it back with 7% annual interest after 10 months. How much interest will she pay?
Answer by Theo(13342)

7% annual interest means the interest rate is compound yearly.
the formula for interest rate being compounded yearly is:
f = p * (1 + r) ^ n
f is the future value
p is the present value
r is the interest rate per year.
n is the number of years.

since 10 months = 10/12 of a year, then the formula becomes:
f = 1200 * (1 + .07) ^ (10/12)
this results in f = 1200 * 1.07 ^ (10/12)
this results in f = 1269.602379
since i = f - p, then i = 69.602379
in this formula, i represents the interest, f represents the future value, p represents the present value.